# OGD4All: A Framework for Accessible Interaction with Geospatial Open Government Data Based on Large Language Models

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I INTRODUCTION
Allowing citizens to retrieve relevant open documents and datasets, which directly answer their specific questions, is a fundamental step towards e-governance [Helbing2023DemocracySociety]. While existing Open Government Data (OGD) repositories, e.g., data.gov, data.europe.eu, and opendata.swiss, provide massive data collections, their access has remained limited by non-semantic search engines, inadequate metadata and documentation, as well as the need for specialized expertise. Often, citizens are forced to manually inspect dozens of datasets in order to verify their relevance.
@@
In recent years, LLMs equipped with appropriate tools have demonstrated increasing proficiency in navigating large data contexts [schick_toolformer_2023, lewis_retrieval-augmented_2020, yao_react_2023] and generating task-appropriate analysis code [gao_pal_2023, yin_natural_2022]. However, when directly prompted to analyze data or provide insights related to public life, current frontier LLMs have been repeatedly observed to produce hallucinated, incorrect information [huang_survey_2025]. Even with Retrieval-Augmented Generation (RAG), the exact source underlying the output produced can be hard to find or verify. Further, platforms like Reddit and Facebook are among the most commonly cited sources of LLMs, according to a 2025 study by Semrush. [^4] https://www.semrush.com/blog/ai-mode-comparison-study/, last accessed: 21 August 2025 This, however, is problematic as these platforms primarily host user-generated content that is rarely subject to editorial oversight or fact-checking.
Ultimately, the challenges of citizens to effectively harness OGD has broader societal implications. While many countries have passed laws ensuring freedom of information, in practice, these rights have little impact if the data is only accessible to a small fraction of citizens. In addition, there are further enabling conditions, such as freedom of the press and political agency [peixoto_uncertain_2013]. Widely accessible tools for extracting insights from open data can foster evidence-based political deliberation, support journalism and accountability, and enable local communities to address public issues collaboratively.
@@ II RELATED WORK
Early work on integrating LLMs with data analysis tasks focused on tabular reasoning and tool-calling. This approach has been explored in design-science research [schelhorn_designing_2024], which combines tabular data analysis and tool calling for an OGD assistant. Recent advances extend these capabilities to geospatial contexts, as seen in GeoGPT [zhang_geogpt_2024], GeoChat [wei_geotool-gpt_2025], GeoTool-GPT [wei_geotool-gpt_2025], and related systems, combining LLMs with GIS tools. It has been repeatedly shown that generating code is a more promising approach than calling tools, notably due to better composability, generality, and representation in LLM training corpora [yang_if_2024, wang_executable_2024]. Some GIS-focused systems harnessing these advantages have been proposed, for example, with LLM-Geo [li_autonomous_2023] or ChatGeoAI [mansourian_chatgeoai_2024].
@@
In parallel to the aforementioned academic works, government initiatives have aimed to improve access to OGD through conversational and visual interfaces [^5] https://stp.wien.gv.at/wienbotwidget/widget/index/, last accessed: 3 September 2025666https://buenosaires.gob.ar/innovacionytransformaciondigital/boti, last accessed: 3 September 2025. However, these applications typically rely on static question-answer pairs or dialogue trees, drastically limiting their ability to handle personalized or context-specific citizen questions. More recent OGD applications like ZüriCityGPT [^7] https://chat.zuericitygpt.ch, last accessed: 9 October 2025 demonstrate growing civic experimentation with LLM-based municipal data assistants. Yet, they often implement a traditional RAG pipeline and thus solely utilize textual documents rather than (geospatial) datasets that require processing, and typically lack quantitative evaluations. Overall, research on OGD highlights the need for frameworks bridging existing concerns while advancing trustworthy and explainable AI in governance [Floridi2019ASociety, Shneiderman2020BridgingPractice].
@@ III-A Dataset Retrieval
A naive retrieval strategy would be to load all metadata documents into an LLM’s context window and instruct it to identify the datasets relevant to a given question. However, this approach does not scale to a large number of documents for LLMs with small context windows. Even for models with large context windows, it incurs high cost and time requirements. Therefore, OGD4All uses a scalable approach where a metadata database is queried, and only the most relevant documents are presented to the LLM. A semantic retrieval strategy embeds both metadata and queries into a dense vector space and matches them using cosine similarity, thereby overcoming the vocabulary problem [furnas_vocabulary_1987] common in keyword search like BM25. For example, a question about “bicycles in the city” can successfully match a document described only with the term “urban cycling”. To this effect, OGD4All utilizes OpenAI’s text-embedding-3-large model [^8] https://platform.openai.com/docs/models/text-embedding-3-large, last accessed 29 July 2025, which further provides multilingual embeddings.
@@ III-B Dataset Analysis
The task of dataset analysis in OGD4All is defined as follows: Given a user’s multimodal question about OGD and a set of datasets required to answer the question and corresponding metadata, generate one or multiple Python code snippets that appropriately process these datasets and produce a textual answer, as well as an interactive map, a plot, and/or a table if applicable. This task definition focuses on the generation of code instead of an appropriate sequence of tool calls (typically in JSON format), as prior work has argued for code-based approaches to be more effective [yang_if_2024, wang_executable_2024]. Furthermore, code is more transparent and explainable than potentially “black-box” tool calls. The programming language Python was chosen for this task as it consistently ranks high in LLM-oriented code generation benchmarks comparing programming languages [cassano_multipl-e_2022]. Further, it is a highly popular choice in both the GIS and Data Analysis domains, with numerous mature and comprehensive tools and libraries available.
The analysis stage is implemented through a coding agent that works towards answering the user’s question step-by-step. It does so through successive iterations of the following sub-steps: formulating a plan for the current step, generating associated code, executing it in a safe, persistent Python environment, and observing the logged output, as proposed in the CodeAct framework [wang_executable_2024]. The coding agent can recover from errors, as in such cases, the error message is provided to it, and it is instructed to generate a corrected code snippet, enabling a limited form of cognitive reflection. This cycle ends when the LLM agent decides to call the final_answer tool, which communicates a list of results to the user, or if the maximum number of steps (20 by default) is reached.
@@ III-C Benchmark
To reproducibly assess the system across key metrics such as correctness, recall, and latency for different LLMs, a dedicated OGD4All benchmark that can be run with various configurations was developed. Our OGD-centric benchmark was created to cover both dataset retrieval and analysis, and it includes tasks with both geospatial and regular tabular operations over potentially multiple datasets. It features 169 manually crafted and verified question-answer pairs, from 70 parameterized templates (e.g., varying a district name or location), each accompanied by a list of relevant datasets and a ground-truth Python script for producing the answer. To enable a robust and repeatable evaluation, the questions are precisely formulated to have an unambiguous and unique answer. The benchmark also includes 30 negative examples, specifically questions where no answer exists based on the available data. This serves to test the system’s ability to reject such inputs. This critical ability is often overlooked in LLM benchmarks, and the commonly employed reinforcement learning based on human feedback has been shown to even worsen it, which leads to LLMs generating hallucinated answers to questions they do not have factual knowledge about [liang_machine_2025].
@@
Table I provides an overview of the utilized metrics. The correctness metric is computed using an LLM-as-a-judge approach [gu_survey_2025] over the 169 non-negative examples. Although reference answers are typically concise and often comprise only a number, name, or location, this method was chosen due to the variability in answer formats and styles generated by LLMs. Preliminary experiments demonstrated that our LLM judge based on GPT-4o is more robust in handling alternate answer representations than a traditional substring match or fuzzy match. Recall, precision, and accuracy are reported as averages over the 199 benchmark questions. In contrast, for latency, token consumption, and API cost, we report the median, as those output distributions were observed to be non-normal, often multimodal, or exhibiting a right skew. Retrieval and analysis performance are assessed separately for interpretability; i.e., the ground-truth relevant datasets are supplied for the analysis benchmark runs.
@@ IV-D Limitations and Future Work
As seen in Figures 5 & 6, current frontier models already saturate our custom benchmark with near-perfect scores in recall, answerability, and correctness. While this limits the discriminability for future frontier models, it also clearly indicates that current models can be used for reliably solving typical, short retrieval and analysis tasks related to OGD. Benchmark complexity could be increased by augmenting the benchmark with question-answer pairs containing domain expert knowledge collected from experts and local knowledge. Furthermore, each benchmark question had a well-defined and unique answer. The system’s behavior for ambiguous or open questions and prompts was also not evaluated. However, in the geospatial context, ambiguity in the formulation of a question is a common issue, as discussed in detail by Mai et al. [mai_geographic_2021].
@@ ACKNOWLEDGMENT
The authors acknowledge that they were able to use the AI infrastructure of Esri to interface with various LLMs. JA and DH are grateful for support by the project “CoCi: Co-Evolving City Life”, which received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 833168.

### v2

@@ I INTRODUCTION
Allowing citizens to retrieve relevant open documents and datasets, which directly answer their specific questions, is a fundamental step towards e-governance [6]. While existing Open Government Data (OGD) repositories, e.g., data.gov, data.europe.eu, and opendata.swiss, provide massive data collections, their access has remained limited by non-semantic search engines, inadequate metadata and documentation, as well as the need for specialized expertise. Often, citizens are forced to manually inspect dozens of datasets in order to verify their relevance.
@@
In recent years, LLMs equipped with appropriate tools have demonstrated increasing proficiency in navigating large data contexts [16, 9, 21] and generating task-appropriate analysis code [4, 22]. However, when directly prompted to analyze data or provide insights related to public life, current frontier LLMs have been repeatedly observed to produce hallucinated, incorrect information [7]. Even with Retrieval-Augmented Generation (RAG), the exact source underlying the output produced can be hard to find or verify. Further, platforms like Reddit and Facebook are among the most commonly cited sources of LLMs, according to a 2025 study by Semrush. [^4] https://www.semrush.com/blog/ai-mode-comparison-study/, last accessed: 21 August 2025 This, however, is problematic as these platforms primarily host user-generated content that is rarely subject to editorial oversight or fact-checking.
Ultimately, the challenges of citizens to effectively harness OGD has broader societal implications. While many countries have passed laws ensuring freedom of information, in practice, these rights have little impact if the data is only accessible to a small fraction of citizens. In addition, there are further enabling conditions, such as freedom of the press and political agency [14]. Widely accessible tools for extracting insights from open data can foster evidence-based political deliberation, support journalism and accountability, and enable local communities to address public issues collaboratively.
@@ II RELATED WORK
Early work on integrating LLMs with data analysis tasks focused on tabular reasoning and tool-calling. This approach has been explored in design-science research [15], which combines tabular data analysis and tool calling for an OGD assistant. Recent advances extend these capabilities to geospatial contexts, as seen in GeoGPT [23], UrbanLLM [8], GeoTool-GPT [19], and related systems, combining LLMs with GIS tools. It has been repeatedly shown that generating code is a more promising approach than calling tools, notably due to better composability, generality, and representation in LLM training corpora [20, 18]. Some GIS-focused systems harnessing these advantages have been proposed, for example, with LLM-Geo [10] or ChatGeoAI [13].
@@
In parallel to the aforementioned academic works, government initiatives have aimed to improve access to OGD through conversational and visual interfaces [^5] https://stp.wien.gv.at/wienbotwidget/widget/index/, last accessed: 3 September 2025, [^6] https://buenosaires.gob.ar/innovacionytransformaciondigital/boti, last accessed: 3 September 2025. However, these applications typically rely on static question-answer pairs or dialogue trees, drastically limiting their ability to handle personalized or context-specific citizen questions. More recent OGD applications like ZüriCityGPT [^7] https://chat.zuericitygpt.ch, last accessed: 9 October 2025 demonstrate growing civic experimentation with LLM-based municipal data assistants. Yet, they often implement a traditional RAG pipeline and thus solely utilize textual documents rather than (geospatial) datasets that require processing, and typically lack quantitative evaluations. Overall, research on OGD highlights the need for frameworks bridging existing concerns while advancing trustworthy and explainable AI in governance [2, 17].
@@ III-A Dataset Retrieval
A naive retrieval strategy would be to load all metadata documents into an LLM’s context window and instruct it to identify the datasets relevant to a given question. However, this approach does not scale to a large number of documents for LLMs with small context windows. Even for models with large context windows, it incurs high cost and time requirements. Therefore, OGD4All uses a scalable approach where a metadata database is queried, and only the most relevant documents are presented to the LLM. A semantic retrieval strategy embeds both metadata and queries into a dense vector space and matches them using cosine similarity, thereby overcoming the vocabulary problem [3] common in keyword search like BM25. For example, a question about “bicycles in the city” can successfully match a document described only with the term “urban cycling”. To this effect, OGD4All utilizes OpenAI’s text-embedding-3-large model [^8] https://platform.openai.com/docs/models/text-embedding-3-large, last accessed 29 July 2025, which further provides multilingual embeddings.
@@ III-B Dataset Analysis
The task of dataset analysis in OGD4All is defined as follows: Given a user’s multimodal question about OGD and a set of datasets required to answer the question and corresponding metadata, generate one or multiple Python code snippets that appropriately process these datasets and produce a textual answer, as well as an interactive map, a plot, and/or a table if applicable. This task definition focuses on the generation of code instead of an appropriate sequence of tool calls (typically in JSON format), as prior work has argued for code-based approaches to be more effective [20, 18]. Furthermore, code is more transparent and explainable than potentially “black-box” tool calls. The programming language Python was chosen for this task as it consistently ranks high in LLM-oriented code generation benchmarks comparing programming languages [1]. Further, it is a highly popular choice in both the GIS and Data Analysis domains, with numerous mature and comprehensive tools and libraries available.
The analysis stage is implemented through a coding agent that works towards answering the user’s question step-by-step. It does so through successive iterations of the following sub-steps: formulating a plan for the current step, generating associated code, executing it in a safe, persistent Python environment, and observing the logged output, as proposed in the CodeAct framework [18]. The coding agent can recover from errors, as in such cases, the error message is provided to it, and it is instructed to generate a corrected code snippet, enabling a limited form of cognitive reflection. This cycle ends when the LLM agent decides to call the final_answer tool, which communicates a list of results to the user, or if the maximum number of steps (20 by default) is reached.
@@ III-C Benchmark
To reproducibly assess the system across key metrics such as correctness, recall, and latency for different LLMs, a dedicated OGD4All benchmark that can be run with various configurations was developed. Our OGD-centric benchmark was created to cover both dataset retrieval and analysis, and it includes tasks with both geospatial and regular tabular operations over potentially multiple datasets. It features 169 manually crafted and verified question-answer pairs, from 70 parameterized templates (e.g., varying a district name or location), each accompanied by a list of relevant datasets and a ground-truth Python script for producing the answer. To enable a robust and repeatable evaluation, the questions are precisely formulated to have an unambiguous and unique answer. The benchmark also includes 30 negative examples, specifically questions where no answer exists based on the available data. This serves to test the system’s ability to reject such inputs. This critical ability is often overlooked in LLM benchmarks, and the commonly employed reinforcement learning based on human feedback has been shown to even worsen it, which leads to LLMs generating hallucinated answers to questions they do not have factual knowledge about [11].
@@
Table I provides an overview of the utilized metrics. The correctness metric is computed using an LLM-as-a-judge approach [5] over the 169 non-negative examples. Although reference answers are typically concise and often comprise only a number, name, or location, this method was chosen due to the variability in answer formats and styles generated by LLMs. Preliminary experiments demonstrated that our LLM judge based on GPT-4o is more robust in handling alternate answer representations than a traditional substring match or fuzzy match. Recall, precision, and accuracy are reported as averages over the 199 benchmark questions. In contrast, for latency, token consumption, and API cost, we report the median, as those output distributions were observed to be non-normal, often multimodal, or exhibiting a right skew. Retrieval and analysis performance are assessed separately for interpretability; i.e., the ground-truth relevant datasets are supplied for the analysis benchmark runs.
@@ IV-D Limitations and Future Work
As seen in Figures 5 & 6, current frontier models already saturate our custom benchmark with near-perfect scores in recall, answerability, and correctness. While this limits the discriminability for future frontier models, it also clearly indicates that current models can be used for reliably solving typical, short retrieval and analysis tasks related to OGD. Benchmark complexity could be increased by augmenting the benchmark with question-answer pairs containing domain expert knowledge collected from experts and local knowledge. Furthermore, each benchmark question had a well-defined and unique answer. The system’s behavior for ambiguous or open questions and prompts was also not evaluated. However, in the geospatial context, ambiguity in the formulation of a question is a common issue, as discussed in detail by Mai et al. [12].
@@ ACKNOWLEDGMENT
The authors acknowledge that they were able to use the AI infrastructure of Esri to interface with various LLMs. JASV and DH are grateful for support by the project “CoCi: Co-Evolving City Life”, which received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 833168.